OpenStudy (anonymous):

Is anyone really good at 8th grade math?

OpenStudy (andrewhaze):

lets see what you got

OpenStudy (anonymous):

me

OpenStudy (anonymous):

I have 5 questions

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

. Two lines, E and F, are represented by the equations given below. Line E: 5x + 5y = 40 Line F: x + y = 8 Which statement is true about the solution to the set of equations? (4 points) It is (8, 40). It is (40, 8). There are infinitely many solutions. There is no solution. 2. The graph shows two lines, K and J. Based on the graph, which statement is correct about the solution to the system of equations for lines K and J? (4 points) (–1, 0) is the solution to line K but not to line J. (0, 4) is the solution to both lines K and J. (1, 1) is the solution to line K but not to line J. (1, 1) is the solution to both lines K and J. 3. Line M is represented by the following equation: x – y = 8 Which equation completes the system that is satisfied by the solution (18, 10)? (4 points) x – y = –28 2x – 2y = 36 x + y = 18 2x – y = 26 4. Two lines, A and B, are represented by the following equations: Line A: y = x – 1 Line B: y = –3x + 11 Which of the following options shows the solution to the system of equations and explains why? (4 points) (3, 2), because the point lies between the two axes (3, 2), because one of the lines passes through this point (3, 2), because the point does not lie on any axis (3, 2), because both lines pass through this point 5. Line H is represented by the following equation: 2x + 2y = 8 What is most likely the equation for line K so the set of equations has infinitely many solutions? (4 points) x + y = 4 x + 2y = 2 2x + y = 6 x – y = 8

OpenStudy (andrewhaze):

ok

OpenStudy (andrewhaze):

where here to help

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

do you guys see them? @andrewhaze

OpenStudy (andrewhaze):

no

OpenStudy (andrewhaze):

imma confused im usually good at graphs :O

OpenStudy (anonymous):

@ilovemath your pic is not right please change it thank you

OpenStudy (ankit042):

It is confusing let us do one question at a time

OpenStudy (ankit042):

for the first question they are basically same line so infinite solutions

OpenStudy (anonymous):

oh okay do you know anyone who could help me? :)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Two lines, E and F, are represented by the equations given below. Line E: 5x + 5y = 40 Line F: x + y = 8 Which statement is true about the solution to the set of equations? (4 points) It is (8, 40). It is (40, 8). There are infinitely many solutions. There is no solution.

OpenStudy (anonymous):

thats #1

OpenStudy (anonymous):

and @illkillyou12 what are you talking about?

OpenStudy (anonymous):

thanks!!!

OpenStudy (anonymous):

The graph shows two lines, K and J. Based on the graph, which statement is correct about the solution to the system of equations for lines K and J? (4 points) (–1, 0) is the solution to line K but not to line J. (0, 4) is the solution to both lines K and J. (1, 1) is the solution to line K but not to line J. (1, 1) is the solution to both lines K and J.

OpenStudy (anonymous):

This stuff is easy. Hold on though I got to write the questions down

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

sorry my computer crashed

OpenStudy (andrewhaze):

their infinite solutions

OpenStudy (andrewhaze):

im gonna show you the anal to @saber2th17 ok @Mlink1290

OpenStudy (andrewhaze):

nvm he's pm you the answers